OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = Cube root of quantity x minus four.

OpenStudy (mertsj):

Did you find f(g(x)) ?

OpenStudy (anonymous):

No?

hero (hero):

Writing it out makes no clear distinction on whether g(x) is \(\sqrt[3]{x-4}\) or \(\sqrt[3]{x} - 4\)

hero (hero):

If you don't know how to use latex, just simply write something like cbrt(x) - 4 or cbrt(x - 4)

OpenStudy (mertsj):

\[f(x)=x^3+4\] \[g(x)=\sqrt[3]{x-4}\]

OpenStudy (mertsj):

\[f(g(x))=(\sqrt[3]{x-4})^3+4=(x-4)+4=x\]

OpenStudy (mertsj):

Do you understand that?

OpenStudy (anonymous):

My bad its cbrt (x-4)

OpenStudy (mertsj):

Do you know that the cube root and the third root is the same thing?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So, therefore, what I wrote is correct.

OpenStudy (anonymous):

Yes no im not saying that it isn't I was replying to Hero. Sorry! And as far as understanding I am pretty sure you plugged in g(x) into f(x) but Im having troubles understanding where you plugged It in

OpenStudy (mertsj):

If a problem says to find the f(3) that means to take the function f and replace every x with a 3. Do you understand that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh Im sorry I see it now! It just takes me a while..

OpenStudy (mertsj):

So if a problem (such as yours) says to find the f(g(x)) that means to take the function f and replace every x with g(x)

OpenStudy (mertsj):

So: